OpenStudy (toxicsugar22):

Find the critical t-value(s) for a one sample t-test given: α = 0.05 df = 13 two-tailed test Select one: A. 2.160 and -2.160 B. 1.960 and -1.960 C. 2.375 and -2.375 D. 1.221 E. 2.639

OpenStudy (anonymous):

sorry I dont know

OpenStudy (johnnewman):

sorry I dont know

OpenStudy (johnnewman):

sorry i cant help

OpenStudy (kropot72):

You need to use a table of the t-distribution for this. Do you have that table?

OpenStudy (kropot72):

In your question, the two-tail probability value is 0.05 and df = 13. So you need to go down the table to the row where df = 13 and across to the column where the two-tail probability value = 0.05. What is the value there?

OpenStudy (kropot72):

Correct. You need to include the negative value of 2.160 in your answer as well as the positive value, the reason being it is a two-tail test.

OpenStudy (kropot72):

If the sample size is n, we have n - 1 degrees of freedom. In your second question, what does df equal?

OpenStudy (kropot72):

But I just explained how you can find the value of df. You just need to subtract 1 from the value given for n. Then you will have the value of df. df = n - 1. df = 17 - 1 = 16.

OpenStudy (kropot72):

Correct. So put a negative sign before the value that you found in the table.

OpenStudy (kropot72):

You are correct, the null hypothesis would be rejected in favor of the alternative hypothesis.

OpenStudy (kropot72):

The p-value is greater than alpha. Therefore we fail to reject the null hypothesis.

OpenStudy (kropot72):

Correct!

OpenStudy (kropot72):

Sorry, I need to log out for a while. Will be back later, however :)

OpenStudy (kropot72):

Later perhaps.

OpenStudy (kropot72):

The test statistic (Z = 2.345) lies outside the rejection region. Therefore we retain the null hypothesis.

OpenStudy (tootzrll):

im sorry i dont know anything about this :[

OpenStudy (musicequalslife):

I believe so... but don't take my word on it because I'm not very positive

OpenStudy (janu16):

2nd question?

OpenStudy (janu16):

i dont know anything about this sorry but let me tag someone who maybe can help @rebeccaxhawaii

OpenStudy (kayders1997):

I have no idea about the null hypothesis I'm sorry :(

OpenStudy (samirmath01):

What is it about?

OpenStudy (mckenna1234):

sorry i don't know any of this @toxicsugar22

OpenStudy (samirmath01):

what do represent \[ \alpha \textrm{ and df } ?\]

OpenStudy (toxicsugar22):

a is allpha and df is degrees of freedom

OpenStudy (toxicsugar22):

@sugarapples @song_of_the_sole @FFASinger @Vestran01 @mayliz @ILovePuppiesLol @mathwizzard3

OpenStudy (toxicsugar22):

@urkhada

OpenStudy (mathwizzard3):

@phi @matlee @mathstudent55

OpenStudy (toxicsugar22):

@DanJS @Directrix

OpenStudy (toxicsugar22):

Investigators are interested in whether a new diet lowers total cholesterol in a group of individuals. They take a sample of 31 participants on the new diet. They know that the mean cholesterol level of the general population (µ) is 200 mg/dL, but they do not have information on the standard deviation of the population. The mean total cholesterol level in their sample of participants is 196 mg/dL and their sample has a standard deviation of 8 mg/dL. Your research question of interest is: Does the population of people on a new diet have a different mean cholesterol level than the general population? You will use the information above for several of the following questions. Which of the following represents the appropriate null hypothesis (H0), given this research question of interest? Select one: A. H0: µ ≠ 200 B. H0: µ = 200 C. H0: µ < 200 D. H0: µ = 196 E. H0: x̄ = 196

OpenStudy (toxicsugar22):

@pphalke

OpenStudy (mathwizzard3):

@ganeshie8 @phi @mathstudent55 @matlee